# Hierarchical LLMs In-the-loop Optimization for Real-time Multi-Robot Target Tracking under Unknown Hazards

### v1

@@ Abstract
In this paper, we propose a hierarchical Large Language Models (LLMs) in-the-loop optimization framework for real-time multi-robot task allocation and target tracking in an unknown hazardous environment subject to sensing and communication attacks. We formulate multi-robot coordination for tracking tasks as a bi-level optimization problem, with LLMs to reason about potential hazards in the environment and the status of the robot team and modify both the inner and outer levels of the optimization. The inner LLM adjusts parameters to prioritize various objectives, including performance, safety, and energy efficiency, while the outer LLM handles online variable completion for team reconfiguration. This hierarchical approach enables real-time adjustments to the robots’ behavior. Additionally, a human supervisor can offer broad guidance and assessments to address unexpected dangers, model mismatches, and performance issues arising from local minima. We validate our proposed framework in both simulation and real-world experiments with comprehensive evaluations, which provide the potential for safe LLM integration for multi-robot problems.
@@ I Introduction
Multi-robot coordination and planning have been extensively employed in different applications like exploration, search and rescue, and target tracking [1, 2, 3, 4, 5]. Coordinating a team of robots in unknown and adversarial environments becomes significantly challenging due to the increased complexity and unpredictability of conditions. Moreover, sensor and communication failures, potentially caused by adversaries, can lead to critical safety issues, including the inability of robots to localize, observe their surroundings, and maintain team connectivity. While existing approaches have focused on designing algorithms to handle modeled adversarial tasks and environments [6, 7, 8, 9], the mismatch between these models and the unexpected dangers encountered in real-world scenarios makes adaptation difficult.
Figure 1: Hardware experiment of multi-drone target tracking with human input. Two drones are tracking two ground robots within an environment that features two sensing danger zones (in red) and one communication danger zone (in blue) as potential hazards. A human supervisor offers guidance to adjust the drones’ behavior to mitigate risks.
@@
Figure 2: The hierarchical LLMs framework for multi-robot tracking. The robots’ inputs and human inputs are processed by two LLMs: a task LLM and an action LLM. The action LLM operates at a higher frequency for reactive performant adaptation, while the task LLM has a lower frequency to provide strategic guidance and reconfiguration to the system. The optimization solver takes the LLMs’ outputs into executable actions, guiding the multi-robot team in real-time.
To address these challenges, we hierarchically integrate LLMs into multiple feedback loops alongside conventional optimizers, leveraging high-level reasoning and decision-making to combine the strengths of both approaches, as illustrated in Fig. 2. We formulate the task assignment and target tracking problem as a bi-level optimization, utilizing the understanding and reasoning capabilities of LLMs for distinct subproblems. Specifically, the outer loop LLM identifies task allocation variables to decide which target to track with reasoning significance, while the inner loop LLM tunes parameters to prioritize objectives towards tracking performance, safety, and energy efficiency to guide the behaviors of robots. In addition to in-the-loop numerical feedback, we also enable human interactions in the system to address abnormal behaviors, remind unexpected danger avoidance, and provide more high-level supervision. We evaluate the proposed framework with several scenarios and deploy it on our customized hardware platforms to validate its real-time robustness and capabilities.
@@ II Related Works
Compared with traditional multi-robot coordination methods, risk-aware and danger-aware strategies have gained significant attention in recent years for their potential to improve resilience and robustness in real-world applications [22]. Hazards can cause temporary or permanent damage to robot sensors, disrupt communication links between robots, and degrade the performance of task execution. Authors in [6] addressed these dangers in multi-robot target tracking and explored chance-constrained optimization that effectively balances minimizing the risk of failures with maximizing target tracking efficiency. Resiliency and adaptive strategies with recovery from sensor and communication failures were further extended in [23, 7]. To migrate the influence of sensing and communication attacks, the work in [24] employed neural models for the formation control of the robot team. Other methods generated active protections by employing jamming techniques to counter eavesdropping [25]. Additionally, active strategies like visual-based anomaly detection [26, 13] helped ensure safe operations under unforeseen environmental conditions.
However, the environments and adversaries considered in these prior studies are ideal compared to real-world scenarios, which involve challenges such as limited heterogeneous resources, dynamically shifting adversarial zones with changing shapes, and the need for real-time adaptation. With the increasing complexity of tasks and the hazards of the environments, the problem formulation might not always be feasible, and optimal solutions cannot be obtained in allowable time with limited computation resources.
Incorporating LLMs into task allocation and motion planning has shown promising results in areas such as reward design [27], generating subgoals [28, 29], and anomaly detection [13]. While these models can offer valuable strategic input, they are generally ineffective and lack guarantees for managing low-level tasks or direct robot actions when used in isolation, as highlighted in previous research [30, 31, 32]. For more complex scenarios with hard requirements, combining LLMs with conventional optimization techniques can enhance both the feasibility and effectiveness of the solutions. LLMs contribute valuable guidance for iterative refinement and problem formulation adjustments [33, 34], while conventional methods offer stronger guarantees for solution quality and constraint handling. This integration motivates our work to explore the potential for more robust and reliable performance.
@@ III-B Task Allocation and Planning as a Bi-level Optimization
We utilize the multi-robot multi-target tracking problem as an instance of multi-robot coordination with task and motion planning to showcase the capabilities of our framework. The unknown hazards in the environments are modeled as danger zones with probabilistic attacks on the robot team. The problem is defined as follows.
Figure 3: Demonstration of multi-robot target tracking with sensing and communication danger zones. The robot team consists of five robots tracking seven targets in an environment with four danger zones. In (a), one robot is attacked while inside a sensing danger zone, and in (b), it successfully recovers from the attack. In (c), a communication attack causes another robot to lose its communication link, and in (d), this robot recovers communication and resumes information sharing with the team.
Definition 1.
(Multi-robot Target Tracking with Danger Zones [6]): We consider M robots equipped with sensors to track N moving targets in hazardous environments. These hazards are categorized into two types: sensing danger zones and communication danger zones. The objective of the multi-target tracking problem is to maximize tracking performance while balancing the deteriorating impacts caused by attacks from these danger zones.
@@
$\vspace{-0.4cm}\begin{split}\min_{\mathcal{A},\mathbf{u}}\ \ &\mathcal{J}(\mathcal{A},\mathbf{u})=\{\mathcal{J}_{1}(\mathcal{A},\mathbf{u}),\cdots,\mathcal{J}_{n}(\mathcal{A},\mathbf{u})\},\\ {s.t.}~{}\ &\mathcal{H}(\mathcal{A},\mathbf{u})=\bm{0},\quad\mathcal{G}(\mathcal{A},\mathbf{u})\preceq\bm{0},\end{split}$ (1)
@@
Instead of deploying every LLM for each individual agent, we adopt a partially centralized optimization, where task LLM guides task assignment, and action LLM assists in tracking processes. The design of the prompts and their corresponding verification methods will be discussed in the following sections.
@@ III-C Task LLM for Reconfiguration
We query task LLM to output the assignment matrix to avoid optimization formulation with matrix-based hard constraints, which is not straightforward to reason. Task assignment in a multi-robot tracking problem requires comprehensive information about robot capabilities, available resources, and the current status of both robots and dynamic tasks (targets). Additionally, historical data is important to evaluate whether a robot’s assignment should be updated. The following information is provided to the task LLM in a descriptive language format: (1) robot and target status: This includes the dynamic states of robots (e.g., position, velocity), target states, and the current status of robots failures by attacks (e.g., whether they are affected by sensing or communication danger zones). (2) environmental status: Known locations of sensing and communication danger zones, along with their impact on the robots. (3) robot capacities: The number of targets a robot can track at once, which may also be dynamically adjusted based on current conditions. (4) historical data: The most recent $k$ results about states of robots and targets, information on previous danger zone attacks, and prior task assignments. This historical context helps the LLM make informed decisions about potential reassignment needs.
@@ III-C1 Prompt Design
The system prompt is designed to feed into LLM by clearly defining its role as an optimization engine tasked with assigning tracking targets to robots. You are an optimizer with the goal of assigning robots to track targets. The user input provides essential real-time and historical information about robots, targets, and the environment. An example is provided:
The recent k results of status and observations are ... The []th information is as follows. Drones are currently at the following positions: Drone [ID] is at [] .... Targets are currently at the following positions: Target [ID] is at [ ] ... The known sensing zones are: Drone [ID] knows the sensor zone [ID] located at ... No attack has been detected. The last trace of the tracking estimation covariances matrix is []. The optimization constraints and request for LLM outputs are described in the form of
Each drone has the ability to track at most [] targets, and each target should be tracked by at least one drone as possible. Please provide the new tracking assignment for each drone in each line with ’’. This structured approach ensures that the LLM has all necessary data, constraints, and output expectations.
@@ III-C2 Output Verification
We employ two types of checkers: the first ensures that the LLM output adheres to the required format with numerical data, while the second verifies whether the output meets the specified hard constraints. The verification function is designed to ensure that the output from the LLM can be used effectively for low-level optimization. It helps to determine whether the current output should be used or skipped, adjusting the prompt for the next query if needed. The assignment matrix is evaluated using with
$\beta(\mathcal{A})=\mathbf{1}^{T}\max\left(\mathbf{0},\mathcal{A}\mathbf{1}-\mathbf{c}\right)+\mathbf{1}^{T}\max\left(\mathbf{0},\mathbf{1}-\mathcal{A}^{T}\mathbf{1}\right)$ (4)
where $\mathbf{c}=(c_{1},\cdots,c_{M})$ is the vector of robot capacities, $\mathbf{1}$ represent a vector of ones. The first term evaluates the robot capacity constraint as satisfying, where $c_{i}$ is the maximum number of targets the robot $i$ can track. The second term qualifies as each target being tracked by at least one robot. If $\beta(\mathcal{A})>0$, then the constraints are not met and we will skip the current output.
@@ III-D1 Prompt Design
1. tracking error computed by the trace of the estimation covariance matrix of the targets 2. control cost computed by the norm of control input 3. slack variables of safety constraints to avoid sensing danger zones 4. slack variables of safety constraints to avoid communication danger zones We use action LLM to evaluate the relative importance of each objective and adjust the parameters $w_{1},w_{2},w_{3},w_{4}$ accordingly. The weight adjustment is a short-horizon decision. Therefore, we provide the LLM with the current status of robots, targets, and environments, with the last weight values. Similarly, the system prompt is designed as:
You are a multiple objective optimizer with the goal of specifying the weights of each objective function. In addition to status, the weight is provided with the specified output format.
@@ III-E Human Input
We consider human reactions to supervise the system to inspect abnormal behaviors, risks, and uncertainty. Humans are generally not able to directly manage massive numerical data in real-time, like the exact positions of all robots, targets, and danger zones, but can have a sense of the actions of robots given some visualizations of data and real-world unmodeled conditions. The human input is optional for the system, and only triggered and added to the user prompt if the human has comments about the current performance. The additional prompt is at the end of the scenario description, and starts with,
In addition, the human supervisor has some input... We list the following types of input (1) performance qualifications: Humans can provide questions or concerns about the general performance. For example, humans can specify one robot to conduct subtracting, like focusing on the track part of the targets. Humans can also raise questions about the performance, reporting the missing tracking of some target, zigzag actions of robots, etc. (2) risk assessment: For environmental hazards, humans can specify some high-level evaluation like to push the robot team to be more conservative, or detailed information like to avoid a particular danger region if the target is very close to this region. (3) abnormal specifies: When deploying the framework to hardware platforms, sometimes there are extra factors that might affect the performance. For example, a robot might get trapped in a local minimum and not move. Then, a human can specify that the robot doesn’t move for a time and what’s the reason for that. In addition, the danger zone we assume might be more severe than its model and thus needs priority adjustment for safety.
@@ IV Results
IV-A Implement Details
We evaluated the capabilities of our framework through numerical simulations in different scenarios. For our experiments, we leveraged both the GPT-3.5-Turbo [36] and GPT-4o [37] models and configured the temperature to 0 to ensure deterministic responses. The maximum token limit is set to $50\times(\bar{c}+2)$ for the GPT-3.5-Turbo model, where $\bar{c}$ is the average capacity of the robots. For the GPT-4o model, we add an extra 200 tokens allowance as it usually generates more explanations in the output.
During the simulations, the inner LLM was updated every 2-3 steps to manage tuned performance, while the outer LLM was refreshed every 8-10 steps to handle higher-level task assignments. The simulation runs for 300 steps. In real-world experiments, the action LLM was updated every 5 seconds, while the task LLM was triggered every 20-30 seconds to assess task allocation and overall performance. The task LLM incorporated the status history from the past 5 steps to provide context for long horizon decision-making. For more scenarios and parameters, please refer to https://github.com/Zhourobotics/hierarchical-llms.git.
IV-B Simulations
IV-B1 Reasoning Ability
We first varied the number of targets and robots to determine whether LLMs can deliver reasonable solutions in the correct format as the task complexity increases. The tests are conducted in an environment with five sensing danger zones and two communication danger zones, using between two to eight robots and two to eight targets. Each robot’s tracking capability is fixed at half the number of targets, ranging from one to four, to ensure a balanced setup.
To study the performance of different LLM models, we carried out experiments with two different models, GPT-4o and GPT-3.5-Turbo. As a result, GPT-4o consistently produced feasible solutions with a 100% success rate, while the success rate of the GPT-3.5-Turbo model is illustrated in Fig. 4. Note that in all of our experiments, the full information of the system is given to the LLMs. This high success rate indicates the capability of LLMs in understanding and reasoning about constraints when global information is available. For the task LLM, we tested scenarios with and without providing examples of prior task assignments. Even in cases where no prior reference is given, the LLM still outputs feasible solutions that satisfy the necessary constraints, demonstrating its adaptability and reliability under increased task complexity. From the results of our experiments, the performance of Task LLM using the GPT-3.5-Turbo model deteriorates when a small number of robots have to track a large number of targets, which makes it more challenging to satisfy the requirements.
(a) Task LLM
(b) Action LLM
Figure 4: Success rate of both the task LLM (a) and the action LLM (b) with the number of robots and targets.
The average token number using these two models is shown in Fig. 5 and Fig. 6. As expected, the token count generally increases with the growing number of robots and targets due to the rising complexity of task assignments as more robots and targets are introduced. The task LLM, in particular, shows a noticeable increase in token usage with higher target numbers. The action LLM doesn’t have an obvious change because even if the prompt token increases, the output maintains a consistent level of information.
(a) Task LLM
(b) Action LLM
Figure 5: Average token count generated by the GPT-3.5-Turbo model for Task LLM (a) and Action LLM (b) with the number of robots and targets.
(a) Task LLM
(b) Action LLM
Figure 6: Average token count generated by the GPT-4o model for Task LLM (a) and Action LLM (b) with the number of robots and targets.
IV-B2 Modular Evaluation

### v2

@@ Abstract
Real-time multi-robot coordination in hazardous and adversarial environments requires fast, reliable adaptation to dynamic threats. While Large Language Models (LLMs) offer strong high-level reasoning capabilities, the lack of safety guarantees limits their direct use in critical decision-making. In this paper, we propose a hierarchical optimization framework that integrates LLMs into the decision loop for multi-robot target tracking in dynamic and hazardous environments. Rather than generating control actions directly, LLMs are used to generate task configuration and adjust parameters in a bi-level task allocation and planning problem. We formulate multi-robot coordination for tracking tasks as a bi-level optimization problem, with LLMs to reason about potential hazards in the environment and the status of the robot team and modify both the inner and outer levels of the optimization. This hierarchical approach enables real-time adjustments to the robots’ behavior. Additionally, a human supervisor can offer broad guidance and assessments to address unexpected dangers, model mismatches, and performance issues arising from local minima. We validate our proposed framework in both simulation and real-world experiments with comprehensive evaluations, demonstrating its effectiveness and showcasing its capability for safe LLM integration for multi-robot systems.
@@ I Introduction
Multi-robot coordination and planning have been extensively employed in different applications like exploration, search and rescue, and target tracking [1, 2, 3, 4, 5]. Coordinating a team of robots in unknown and adversarial environments becomes significantly challenging due to the increased complexity and unpredictability of conditions. Moreover, sensor and communication failures, potentially caused by adversaries, can lead to critical safety issues, including the inability of robots to localize, observe their surroundings, and maintain team connectivity. While existing approaches have focused on designing algorithms to handle modeled adversarial tasks and environments [6, 7, 8], the mismatch between these models and the unexpected dangers encountered in real-world scenarios makes adaptation difficult.
Figure 1: Real-world experiments. Two drones track ground robots in an environment with two sensing zones (red) and one communication zone (blue), guided by human supervision to mitigate risks.
@@
Figure 2: The hierarchical LLM framework for multi-robot tracking processes both robot and human inputs through two LLMs: a high-frequency action LLM for reactive performance adaptation, and a low-frequency task LLM to provide strategic guidance and reconfiguration to the system. The optimization solver takes the LLMs’ outputs into executable actions, guiding the multi-robot team in real-time.
To address these challenges, we propose a hierarchical framework that combines the reasoning capabilities of LLMs with optimizers, enabling real-time, robust multi-robot coordination in hazardous environments. We formulate the task allocation and planning for target tracking under communication and sensing attacks as a bi-level optimization, and employ a hierarchical LLM architecture to handle reasoning tasks at different levels. The outer loop LLM identifies task allocation variables to decide which target to track with reasoning significance, while the inner loop LLM adjusts parameters to balance performance, safety, and energy efficiency during execution. In addition to numerical feedback, we also enable human interactions in the system to address abnormal behaviors, remind of unexpected danger avoidance, and provide more high-level supervision. We validate the proposed framework through diverse simulated scenarios and real-world hardware experiments (Fig. 1), demonstrating its real-time robustness and adaptability. The contributions of the paper are as follows:
  • A novel bi-level optimization framework that incorporates LLMs for real-time, language-guided task allocation and planning in adversarial environments.
  • A hierarchical system design that integrates LLM reasoning with optimizers, supporting constraint enforcement and automatic output validation.
  • Human interaction for high-level supervision and risk mitigation, validated through comprehensive simulation and hardware experiments.
@@ II Related Works
Compared with traditional multi-robot coordination methods, risk-aware and danger-aware strategies have gained significant attention in recent years for their potential to improve resilience and robustness in real-world applications [21]. Hazards can cause temporary or permanent damage to robot sensors, disrupt communication links between robots, and degrade the performance of task execution. Authors in [6] addressed these dangers in multi-robot target tracking and explored chance-constrained optimization that effectively balances minimizing the risk of failures with maximizing target tracking efficiency. Resiliency and adaptive strategies with recovery from sensor and communication failures were further extended in [22]. To mitigate the influence of sensing and communication attacks, the work in [23] employed neural models for the formation control of the robot team. Other methods generated active protections by employing jamming techniques to counter eavesdropping [24]. Additionally, active strategies like visual-based anomaly detection [25, 12] helped ensure safe operations under unforeseen environmental conditions.
However, the environments and adversaries considered in these prior studies are ideal compared to real-world scenarios, which involve challenges such as limited heterogeneous resources, dynamically shifting adversarial zones with changing shapes, and the need for real-time adaptation. With the increasing complexity of tasks and the hazards of the environments, the problem formulation might not always be feasible, and optimal solutions cannot be obtained in an allowable time with limited computation resources.
Incorporating LLMs into task allocation and motion planning has shown promising results in areas such as reward design [26], generating subgoals [27, 28], and anomaly detection [12]. While these models can offer valuable strategic input, they are generally ineffective and lack safety guarantees for managing low-level tasks or direct robot actions when used in isolation, as highlighted in previous research [29, 30, 31, 32]. Recent work has begun to try to safety constraints into LLM-based task planning [33, 34]. Furthermore, combining LLMs with conventional optimization techniques can improve both the feasibility and effectiveness of the solutions. LLMs can contribute valuable guidance for iterative refinement and problem formulation adjustments [35, 36], while conventional methods offer stronger guarantees for solution quality and constraint handling. This integration motivates our work to explore the potential for more robust and reliable performance in multi-agent coordination.
@@ III-A System Overview
Figure 3: Demonstration of multi-robot target tracking with sensing and communication danger zones. The robot team consists of five robots tracking seven targets in an environment with four danger zones. (a–b) show a robot entering a sensing zone, being attacked, and recovering. (c–d) illustrate a communication failure and subsequent recovery.
@@ III-B Task Allocation and Planning as a Bi-level Optimization
We utilize the multi-robot multi-target tracking problem as an instance of multi-robot coordination with task and motion planning to showcase the capabilities of our framework. The unknown hazards in the environment are modeled as danger zones with probabilistic attacks on the robot team. Following the formulation in [5, 6], a team of $M$ sensor-equipped robots is assigned to track $N$ moving targets within such hazardous environments. These hazards are categorized into two types: sensing danger zones and communication danger zones. The objective of the multi-target tracking problem is to maximize tracking performance while balancing the deterioration caused by attacks from these danger zones.
@@
$\begin{split}\min_{\mathcal{A},\mathbf{u}}\ \ &\mathcal{J}(\mathcal{A},\mathbf{u})=\{\mathcal{J}_{1}(\mathcal{A},\mathbf{u}),\cdots,\mathcal{J}_{n}(\mathcal{A},\mathbf{u})\},\\ {s.t.}~\ &\mathcal{H}(\mathcal{A},\mathbf{u})=\bm{0},\quad\mathcal{G}(\mathcal{A},\mathbf{u})\preceq\bm{0},\end{split}$ (1)
@@
Instead of deploying an LLM for each individual agent, we adopt a partially centralized optimization approach, solved with solvers such as FORCES Pro [37]. In this framework, a task LLM guides task assignment, while an action LLM assists in tracking processes. The prompt designs and verification methods will be discussed in the following sections.
@@ III-C Task LLM for Reconfiguration
We query the task LLM to output the assignment matrix to avoid optimization formulation with matrix-based hard constraints, which is not straightforward to reason. Task assignment in a multi-robot tracking problem requires comprehensive information about robot capabilities, available resources, and the current status of both robots and dynamic tasks (targets). Additionally, historical data is important to evaluate whether a robot’s assignment should be updated. The following information is provided to the task LLM in a descriptive language format: (1) robot and target status: This includes the dynamic states of robots (e.g., position, velocity), target states, and the current status of robots failures by attacks (e.g., whether they are affected by sensing or communication danger zones). (2) environmental status: Known locations of sensing and communication danger zones, along with their impact on the robots. (3) robot capacities: The number of targets a robot can track at once, which may also be dynamically adjusted based on current conditions. (4) historical data: The most recent $k$ results about states of robots and targets, information on previous danger zone attacks, and prior task assignments. This historical context helps the LLM make informed decisions about potential reassignment needs.
@@ III-C1 Prompt Design
The system prompt is designed to feed into LLM by clearly defining its role as an optimization engine tasked with assigning tracking targets to robots.
You are an optimizer with the goal of assigning robots to track targets.
The user input provides essential real-time and historical information about robots, targets, and the environment. An example is provided:
The recent k results of status and observations are ... The []th information is as follows. Drones are currently at the following positions: Drone [ID] is at [] .... Targets are currently at the following positions: Target [ID] is at [] ... The known sensing zones are: Drone [ID] knows the sensor zone [ID] located at ... No attack has been detected. The last trace of the tracking estimation covariance matrix is [].
The optimization constraints and request for LLM outputs are described in the form of
Each drone has the ability to track at most [] targets, and each target should be tracked by at least one drone as possible. Please provide the new tracking assignment for each drone in each line with ’’.
This structured approach ensures that the LLM has all the necessary data, constraints, and output expectations.
@@ III-C2 Output Verification
We employ two types of checkers: the first ensures that the LLM output adheres to the required format with numerical data, while the second verifies whether the output meets the specified hard constraints. The verification function is designed to ensure that the output from the LLM can be used effectively for low-level optimization. It helps to determine whether the current output should be used or skipped, adjusting the prompt for the next query if needed. The assignment matrix is evaluated using
$\beta(\mathcal{A})=\mathbf{1}^{T}\max(\mathbf{0},\mathcal{A}\mathbf{1}-\mathbf{c})+\mathbf{1}^{T}\max(\mathbf{0},\mathbf{1}-\mathcal{A}^{T}\mathbf{1}),$ (4)
where $\mathbf{c}=(c_{1},\cdots,c_{M})$ is the vector of robot capacities, $\mathbf{1}$ represent a vector of ones. The first term evaluates the robot capacity constraint as satisfying, where $c_{i}$ is the maximum number of targets the robot $i$ can track. The second term qualifies as each target being tracked by at least one robot. If $\beta(\mathcal{A})>0$, then the constraints are not met, and we will skip the current output.
@@ III-D1 Prompt Design
1. tracking error computed by the trace of the estimation covariance matrix of the targets 2. control cost computed by the norm of control input 3. slack variables of safety constraints to avoid sensing danger zones 4. slack variables of safety constraints to avoid communication danger zones
We use action LLM to evaluate the relative importance of each objective and adjust the parameters $w_{1},w_{2},w_{3},w_{4}$ accordingly. The weight adjustment is a short-horizon decision. Therefore, we provide the LLM with the current status of robots, targets, and environments, with the last weight values. Similarly, the system prompt is designed as:
You are a multiple objective optimizer with the goal of specifying the weights of each objective function.
In addition to status, the weight is provided with the specified output format.
@@ III-E Human Input
We consider human reactions to supervise the system to inspect abnormal behaviors, risks, and uncertainty. Though humans are generally not able to directly manage massive numerical data in real-time, like the exact positions of all robots, targets, and danger zones, but can have a sense of the actions of robots given some visualizations of data and real-world unmodeled conditions. Human input is optional and added to the user prompt if the human has comments about the current performance.
We list the following types of input (1) performance qualifications: Humans can provide questions or concerns about the general performance. For example, humans can specify one robot to conduct subtraction, like focusing on the track part of the targets. Humans can also raise questions about the performance, reporting the missing tracking of some targets, zigzag actions of robots, etc. (2) risk assessment: For environmental hazards, humans can specify some high-level evaluation, such as pushing the robot team to be more conservative, or detailed information, such as avoiding a particular danger region if the target is very close to this region. (3) abnormal specifies: When deploying the framework to hardware platforms, sometimes there are extra factors that might affect the performance. For example, a robot might get trapped in a local minimum and not move. Then, a human can specify that the robot doesn’t move for a time and what’s the reason for that. In addition, the danger zone we assume might be more severe than its model and thus needs priority adjustment for safety.
@@ IV Results
This section presents a series of simulations and real-world experiments designed to evaluate the following hypotheses:
  • H1.
    LLMs are able to reconfigure the task allocation of robots based on the structured language feedback of robots and environment information.
  • H2.
    LLMs are able to adjust the objective weights based on the text feedback.
  • H3.
    Integrating LLMs into the system can boost tracking task performance and reduce risks.
  • H4.
    The proposed LLM-in-the-loop framework is deployable in real-time in robotic hardware.
IV-A Simulation Evaluation
IV-A1 Implement Details
We evaluated the capabilities of our framework through numerical simulations in different scenarios. For our experiments, we utilized GPT-4o [38], GPT-4.1-mini, LLaMA-3-70B [39], and DeepSeek-V3 [40], setting the temperature to 0 to ensure deterministic outputs. The token limit is $50\times(\bar{c}+2)$, where $\bar{c}$ is the average robot tracking capacity. GPT-4o gets extra tokens for its detailed outputs. During the simulations, the inner LLM was updated every 2-3 steps to manage tuned performance, while the outer LLM was refreshed every 8-10 steps to handle higher-level task assignments. The simulation runs for 300 steps.
Figure 4: Success rate and token count generated by different models for the Task LLM (a) and Action LLM (b), evaluated under varying task loads and environmental complexities.
IV-A2 Reasoning Ability
We design scenarios with varying levels of complexity to evaluate the performance of our framework, focusing on two key dimensions: task load, quantified by the robot-to-target ratio, and environmental complexity, measured by the number and distribution of danger zones. The task load is categorized into three levels as underloaded (0.5), balanced (1.0), and overloaded (2.0) based on the ratio of targets to robots. For example, with four robots, these levels correspond to tracking 2, 4, and 8 targets, respectively, assuming each robot can track up to 3 targets. Environmental complexity is similarly categorized into three levels, with 2, 4, and 8 danger zones representing low, medium, and high complexity, respectively. This design allows us to isolate and analyze the impact of internal resource limits and external environmental stressors on tracking performance, robustness, and real-time feasibility.
We evaluated four LLMs across the different scenarios, with Fig. 4 summarizing success rates and token usage for both the Task and Action LLMs. All models achieved a 100% success rate for the Action LLM, For the Task LLM, success rates are shown in the first row of the figure. The token count for the Task LLM generally increases with the number of robots and targets, reflecting the growing complexity of task assignments. In contrast, the Action LLM’s token usage remains relatively stable, as the output complexity does not scale significantly with the prompt size. We also tested the Task LLM with and without prior examples of task assignments. Even without such references, the models consistently generated feasible outputs that satisfied all constraints, demonstrating strong generalization and adaptability under increasing task complexity. These results demonstrate that LLMs can effectively generate task assignments (H1) and dynamically adjust objective weights (H2) based on structured language inputs.
IV-A3 Ablation Study